Head Green Keepers Report December 2017

Current Cutting Heights

Greens                4.7mm

Tees/Collars     10mm

Fairways            20mm

Semi Rough     30mm

Rough                 50mm

Course Maintenance

Weekends rainfall

Cutting has been limited more recently after some prolonged and heavy rainfall. Over the evening of Saturday 9th December some 28mm of rain occurred leaving the course unplayable on the following morning.

 

 

 

 

 

 

Slitting greens 

Slit tine aeration is being undertaken on a monthly basis to help with root development and to keep the greens root zones draining. Light liquid feeding is taking place when conditions are conducive to spraying. This is a combination of iron, potassium and a very negligible amount of nitrogen. This helps stave off disease and control cast worms.

 

 

 

 

 

 

 

Worms

Worm control is an issue under hot debate within the green keeping industry. Our last tried and tested method of control, carbendazim has been removed from use and now we face a challenge of a new magnitude in an effort to suppress casting. Worm casts make mowing difficult and are smeared by the cutters rollers in addition to golfers, trolleys etc. thus impeding drainage. Weed seeds are brought to the surface and weed establishment (especially daises) is encouraged by the fertility of the soil. Sand dressing and ph control are the two actions to improve turf playability followed by selective weed control when strong growth recommences.

Bunkers

Sand (Newton Abbot) is still to be installed into the bunkers on holes 2 and 14. We are hopeful of getting these completed prior to the Christmas period.

Practice Ground Mat

The installation of an 8x4m practice mat on the top tier of the practice tee is underway and we are planning completion by this coming weekend. A short time of settlement will be required before it is brought into use (possibly 2 weeks)

This facility will allow resting of the lower grass tee and an opportunity to improve its condition.

Practice tee installation

 

 

 

 

 

 

 

 

 

Other Items of Interest

The hollow located within the 16th green has now been levelled. We will be monitoring this over the coming season to see if any further settlement should occur.

Down time is being used to construct and paint new fixed marker points to replace existing ones that have lost their colour.

The new top dresser and trailer have now been delivered. Both will be invaluable additions over the coming months.

 

 

 

 

 

16th Green                                              New Markers                                             Top Dresser